# STORMY : A Real-time Triggering Framework using Yamagawa Solar Spectrograph for Active Solar Emission Observations with the MWA

### v1

@@ Abstract
Some of the most interesting insights into solar physics and space weather come from studying radio emissions associated with solar activity, which remain inherently unpredictable. Hence, a real-time triggering system is needed for solar observations with the versatile new-generation radio telescopes to efficiently capture these episodes of solar activity with the precious and limited solar observing time. We have developed such a system, Solar Triggered Observations of Radio bursts using MWA and Yamagawa (STROMY) for the Murchison Widefield Array (MWA), the precursor for the low frequency telescope of upcoming Square Kilometre Array Observatory (SKAO). It is based on near-real-time data from the Yamagawa solar spectrograph, located at a similar longitude to the MWA. We have devised, implemented, and tested algorithms to perform an effective denoising of the data to identify signatures of solar activity in the Yamagawa data in near real-time. End-to-end tests of triggered observations have been successfully carried out at the MWA. STORMY is operational at the MWA for the routine solar observations, a timely development in the view of the ongoing solar maximum. We present this new observing framework and discuss how it can enable efficient capturing of event-rich solar data with existing instruments, like the LOw Frequency ARray (LOFAR), Owens Valley Radio Observatory - Long Wavelength Array (OVRO-LWA) etc., and pave the way for triggered observing with the SKAO, especially the SKA-Low.
@@ 1 Introduction
The observed solar radio emission is a superposition of emissions from a large variety of mechanisms – from incoherent mechanisms like thermal bremsstrahlung and gyromagnetic emission (white1997radio; nindos2020incoherent, e.g.,) to coherent emission mechanisms like plasma emission (ginzburg1959mechanisms; melrose1980emission, e.g.) and electron-cyclotron maser emission (winglee1986electron; tang2013electron, e.g.). The coherent emission mechanisms are capable of producing much higher brightness temperatures and hence give rise to a large increase in the observed intensity. These bright emissions, known as solar radio bursts, are linked with various extreme solar phenomena ranging from energetic CMEs to flares spanning a large range of energies. The radio bursts originate from plasma emission and their spectro-temporal variation in dynamic spectra can show markedly different structures due to different origins of the energetic electrons producing these bursts. Based on their morphology in the dynamic spectrum, they are classified into several types wild; mclean1985solar. Detecting and studying such solar bursts is of utmost importance in understanding their emission mechanisms bastian1998radio; melrose2017coherent, probing the turbulent magnetized coronal medium through which these emissions propagate as well as helping in space-weather research Zhang2014.
@@
Owing to their high brightness, many fine details of the emission can be captured by comparatively simpler and inexpensive single dish/element instrumentation. Technological ease and the comparative simplicity of analysis have implied that non-imaging spectrograph observations have dominated the studies of the Sun by a very wide margin. Despite their obvious limitations, we have learnt a lot from spectroscopic observations (maxwell_swarup; classen2002association; mann_formation; gopalswamy2005type, e.g.,). However, due to the large and potentially simultaneous variations in temporal and spectral structures, brightness temperatures and polarization properties of the solar radio bursts, spectro-polarimetric snapshot imaging with high dynamic range is needed to study these events in their full glory. Generations of former and presently operational radioheliographs like Culgoora Wild1967, Clark Lake Gregley1985 NoRH nakajima; takano2007upgrade; shibasaki2013long, the Nançay Radioheliograph kerdraon; saint2012decade and Gauribidnur Ramesh1998 etc. enabled routine imaging of the Sun at a few discrete frequencies, helping in the study of solar flares and noise storms. Despite this, the instrumentation, algorithmic and computational challenges posed by the needs of solar radio imaging have held back progress. More recent solar-dedicated instruments such as the Siberan Radioheliograph srh_2020, Mingantu Spectral Redioheliograph yan2021mingantu, Expanded Owens Valley Solar Array (gary2018microwave, EOVSA;), Owens Valley Radio Observatory - Long Wavelength Array (chhabra2021imaging, OVRO-LWA;), and Daocheng Solar Radio Telescope (Yan2023-DSRT, DSRT;), have made progress in addressing some of the limitations of earlier instruments (gary2023new, see review by). Alongside, over the last decade or so, a new generation of radio interferometers e.g. the LOFAR lofar_paper, the MWA tingay2013murchison, MeerKAT meerkat_2016. In terms of their raw capabilities, these instruments tend to be best-in-class along some/ many of the parameters of interest (e.g. uv-coverage, resolution, spectroscopic imaging capability, calibration fidelity, etc.) and are much better suited to the demanding requirements of solar imaging have become available. These sophisticated instruments, all precursors or pathfinders for the Square Kilometre Array Observatory (SKAO), are designed to be versatile, capable of addressing a vast variety of science areas and optimized for synthesis observations of faint sources. More often than not, considerable effort is required to adapt these instruments and the analysis procedures for solar observations. This effort has and continues to deliver considerable science returns (morosan_2014; reid_kontar; kansabanik2024spectropolarimetric; oberoi2023preparing; shilpi_type2; kansabanik2025solar; dey2025first, e.g.,).
@@
Being verstile and state-of-the-art, the observing time at these new generation facilities is highly oversubsribed. On the other hand, the solar radio bursts are infrequent, transient, and most importantly, unpredictable.
@@
Hence, blind scheduled Solar observations are not only an extremely inefficient way to capture these bursts in large numbers, they also lead to a high likelihood of missing big solar events. One way to overcome this observational limitation is by optimally utilizing the available observing time using a robust and reliable automated near-real-time triggered observation framework. By enabling one to use precious observing time only when some solar activity is known to have just taken place, such a system can vastly increase the efficiency of the limited available observing time to capture instances of solar activity. Having such systems will not only help in boosting efficiency of Solar research with the existing instruments but also will pave a way to prepare for optimized solar observations the upcoming SKAO.
@@
In radio astronomy, triggered observations have been widely used to capture astrophysical transient activities such as FRBs, GRB afterglows, or pulsar glitches where the response times of seconds to hours are acceptable. Solar radio burst triggers needs near-real time response and low-latency setups to configure the interferometer to observe the Sun as soon as the trigger is received. We have developed such a framework for the low-frequency precursor of the SKAO, the MWA, based on near-real-time data from the Yamagawa solar radio spectrograph Iwai_2017. We have named this framework Solar Triggered Observations of Radio bursts using MWA and Yamagawa (STORMY).
@@
The paper is organised in the following way: an overview of the pipeline is given in section 2, section 3 describes the preprocessing of the data, section 4 talks about the event detection and trigger generation, we show the performance test of STORMY in section 6, in section 7 we present a few events captured by our pipeline, we discuss the application of this algorithm to other facilities in section 8.
@@ 2 Overview of Real-time Burst Detection and Triggering Pipeline
While there have been some attempts to develop automated algorithms for the detection of solar radio bursts from dynamic spectrum, they tend to be focused on the detection and characterization of the type of bursts (e.g. type IIs or IIIs). Earlier works by Lobzinn and bonnin2011automated proposed the use of the Hough transform hough1962 on the dynamic spectra. Dayal have developed an automated burst detection pipeline using e-CALLISTO benz2005callisto; benz2009world data, and salmane have developed a Constant False Alarm Rate (CFAR) based technique for automatic detection of radio bursts in data from the Nançay Decametric Array boischot1980new. In recent times, scientists have also successfully implemented different convolutional neural network (CNN) based deep learning models to detect and categorize solar radio bursts (scully2021type; orue2023automatic; deng2024real, e.g.,). However, these algorithms were developed for searching and classifying radio bursts present in the archival data, where the entire morphological information of the burst is available in the dynamic spectrum. In case of generating a trigger for scheduling an interferometric observation, one does not have the luxury of allowing the full solar radio event to play out before raising the trigger. This includes reliably rejecting any radio frequency interference (RFI) that is inevitably present. Developing such a framework to conduct robust and efficient triggered observations requires the following capabilities: (1) near-real-time access to appropriate observations from solar-dedicated telescopes, based on which a trigger can be generated; (2) algorithms and software tools needed to ingest these data and process them to provide a trigger; (3) the framework required for this trigger to be communicated to the telescope under consideration and, (4) scheduling the observations and the acquisition of data. Here we focus on the first two of the aspects mentioned here, and the latter two have already been implemented at the MWA for other astronomical observations and have been adapted for the present purpose.
@@ 2.1 Yamagawa and MWA
Figure 1: This figure shows an example of the real-time update of data from Yamagawa spectrograph and the MWA voltage buffer. For example: the chunk a is available at 2:46 and the next chunk b is available at 2:48, as seen in the figure. When a trigger is sent, the buffer data (red block) is saved and a regular correlator observation (blue arrow) is triggered. We lose about 1-minute of data even after recovering 160s of data from the MWA buffer.
@@
Our solar observing trigger is derived from the data received from the Yamagawa heliospectrograph, located in Japan at a latitude and longitude of 31.204N and 130.617E, respectively. It observes the Sun daily from rise to set in the frequency band from 70 MHz to 9 GHz. Since Yamagawa is at a similar longitude as the MWA (26.703S, 116.671E) and the MWA only observes the Sun above 30 elevation, Yamagawa has complete overlap with the solar observing window of the MWA. We have optimized the Yamagawa system to enable a custom near-real-time datastream. This datastream provides data at 2-minute intervals with a latency of 2 minutes (Fig. 1) and is better suited to our needs than the publicly available data. The data have a spectral and temporal resolution of 1 MHz and 1 second, respectively.
The MWA observes in the 80–300 MHz band and offers 30.72 MHz of instantaneous processed bandwidth. This instantaneous bandwidth is available in 24 chunks of 1.28 MHz each, which can be distributed flexibly across the entire observing frequency range. Solar observations are often carried in the so called picket fence mode where the observing bandwidth is distributed in 1.28 MHz chunks across the full 80–300 MHz bans in a log-spaced manner.
The MWA also offers a unique ability to store the voltage data for each of the polarizations from each of the tiles in a ring buffer morrison2023mwax. These data are extremely voluminous ($\sim$56 TB/hour for 128 tiles) and with 128 tiles the ring buffer can store upto 160s160\ s of raw voltage data. The voltage buffer offers two key advantages – first, it provides access to the data for up to 160s160\ s prior to the trigger arriving at the MWA and thus allowing us to recover much of the event data lost due to the latency in raising the trigger; and second, it provides flexibility in the time and frequency resolution with which to compute correlations (visibilities) down to the Nyquist sampling rate.
On command, this buffer data can be dumped to disk and can later be passed through an offline correlator to obtain visibility data and reduce data volume. The buffer is filled by the datastream from the pointing of the last observation. Hence, we configure the MWA to point at the Sun in the picket fence mode for all of the unallocated daytime hours, when the Sun is above the elevation limit of the MWA (30). So even though no visibility data is recorded, this ensures that the buffer is populated with solar data. Figure 1 shows how the near real-time datastream comes from the Yamagawa spectrograph, the time taken raise the trigger and the start of observations at the MWA. It is clear from the diagram that due to the buffer available at the MWA (red block), we lose a very small amount of data (blue block) due to the latency of the Yamagawa data.
@@ 2.2 Structure of the Pipeline
Figure 2 shows the complete pipeline. After the near-real-time data is downloaded, the data flows through the following parts:
@@ 3.1 Bandpass Correction
The bandshape of an instrument is its spectral response and can vary significantly across the observing band. The observed intensity of the signal that is recorded by each spectral channel gets modified by the instrumental bandshape. Hence, we need to correct for the power levels of each spectral channel by the instrumental bandshape. Although the same-day data can also be used for band shape calculation, identifying quiet time without any solar bursts in the real-time data becomes computationally inefficient, and also, if a very long duration solar signal is present in the data, the median might get affected. Hence, to calculate the bandshape of the Yamagawa spectrograph, we use previous full-day data ($\sim$8 hours). This data is split into chunks each spanning 180 seconds. Now, we calculate the bandshape of each chunk ($B_{i}(\nu)$) by taking median along the temporal axis and then we calculate the final bandshape by taking a median of all the chunks ($B(\nu)=Median[B_{i}(\nu)]$). This median bandshape is used to normalise the dynamic spectra as follows, $S(\nu,t)=\frac{I(\nu,t)}{B(\nu)}$, where, $I(\nu,t)$ denoted raw dynamic spectrum. This normalisation removes instrumental spectral response and solar features become more easily detectable. Figure 3 shows a sample raw data and the previous day bandshape used for bandshape correction. We will use this data throughout this paper to show further steps.
Figure 4: This figure shows the bandpass normalised dynamic spectrum where the varying signals are easily detectable. The yellow rectangles show a periodic RFI that is seen every 300 seconds. The red arrows show narrow band persistent RFIs.
@@ 3.2 RFI Mitigation
After the bandpass normalisation, different spectrally and temporally varying features can be seen with enhanced contrast. However, the dynamic specrtum has different types of unwanted contamination from RFI which can sometimes mimic the true solar signal. We need to remove them before starting to detect the solar signal. Otherwise, we may get false positive detections. Analysing the archival Yamagawa data, we identified a few different types of RFI that can be present in the data (Figure 4) – (1) periodic RFI, (2) narrow band and persistent RFI and (3) transient RFI.
First, we remove the periodic RFI. They are present every day and have a periodicity of five minutes with a duty-cycle of $\sim$10-seconds, but are fragmented across the channels (Refer to the bottom left panel of figure 5). Although they are fragmented, they appear exactly after 300 seconds in each channel (top left panel of figure 5. Let us denote $S_{\nu}(t)$ as time series a channel. We use the find_peaks module from scipy to find peaks with periodicity of 300 seconds by using the following parameters:
@@
Figure 5: This figure shows how the periodic RFI is cleaned. The top left panel shows the dynamic spectrum with fragmented periodic RFI. The top right panel shows the dynamic spectrum after cleaning these RFI. The part within the 2. The x-axes are shared across subplots in each column.
Next, we clean any persistent narrow-band RFI. They span only a few channels in frequency, deviating largely from the background and persistently present for a long time. Since solar emissions can also have significant deviations from the background, extra care should be taken while filtering these RFI. We used a median filter with a sufficiently small window (5 MHz) to arrive at a smoothed representation of the spectrum at a timestamp, $t_{0}$. This is denoted by $S_{\mathrm{m}}(\nu,t=t_{0})$. We calculate the modified $z$-score of the quantity $\delta(\nu,t=t_{0})=[S(\nu,t=t_{0})-S_{m}(\nu,t=t_{0})]$. Compared to regular z-score calculation, we use the modified z-score since it uses median absolute deviation (MAD) and is less affected by extreme outliers. In our case it is calculated as follows:
@@
$S^{\prime}(\nu,t)=\begin{cases}S(\nu,t),&\text{if }z<10\\ S_{m}(\nu,t),&\text{otherwise}\end{cases}$ (3)
@@
Figure 6: The top left and right panels show the dynamic spectrum before and after narrow band RFI removal, respectively. The bottom left and right panels show the spectra of a time slice, as shown by the red vertical line on the dynamic spectrum in the top panel. The red trace in bottom left panel shows the median smoothed spectra ($S_{m}(\nu)$) of the same.

### v2

@@ Abstract
Some of the most interesting insights into solar physics and space weather come from studying radio emissions associated with solar activity, which remain inherently unpredictable. Hence, a real-time triggering system is needed for solar observations with the versatile new-generation radio telescopes to efficiently capture these episodes of solar activity with the precious and limited solar observing time. We have developed such a system, Solar Triggered Observations of Radio bursts using MWA and Yamagawa (STORMY) for the Murchison Widefield Array (MWA), the precursor for the low frequency telescope of upcoming Square Kilometre Array Observatory (SKAO). It is based on near-real-time data from the Yamagawa solar spectrograph, located at a similar longitude to the MWA. We have devised, implemented, and tested algorithms to perform an effective denoising of the data to identify signatures of solar activity in the Yamagawa data in near real-time. End-to-end tests of triggered observations have been successfully carried out at the MWA. STORMY is operational at the MWA for the routine solar observations, a timely development in the view of the ongoing solar maximum. We present this new observing framework and discuss how it can enable efficient capturing of event-rich solar data with existing instruments, like the LOw Frequency ARray (LOFAR), Owens Valley Radio Observatory - Long Wavelength Array (OVRO-LWA) etc., and pave the way for triggered observing with the SKAO, especially the SKA-Low.
@@ 1 Introduction
The observed solar radio emission is a superposition of emissions from a large variety of mechanisms – from incoherent mechanisms like thermal bremsstrahlung and gyromagnetic emission (white1997radio; nindos2020incoherent, e.g.,) to coherent emission mechanisms like plasma emission (ginzburg1959mechanisms; melrose1980emission, e.g.) and electron-cyclotron maser emission (winglee1986electron; tang2013electron, e.g.). The coherent emission mechanisms are capable of producing much higher brightness temperatures and hence give rise to a large increase in the observed intensity. These bright emissions, known as solar radio bursts, are linked with various extreme solar phenomena ranging from energetic CMEs to flares spanning a large range of energies. The radio bursts originate from plasma emission and their spectro-temporal variation in dynamic spectra can show markedly different structures due to different origins of the energetic electrons producing these bursts. Based on their morphology in the dynamic spectrum, they are classified into several types wild; mclean1985solar. Detecting and studying such solar bursts is of utmost importance for understanding their emission mechanisms bastian1998radio; melrose2017coherent, probing the turbulent magnetized coronal medium and studying the space-weather Zhang2014.
@@
Owing to their high brightness, many fine details of the emission can be captured by comparatively simpler and inexpensive single dish/element instrumentation. Technological ease and the comparative simplicity of analysis have implied that non-imaging spectrograph observations have dominated the studies of the Sun by a very wide margin. Despite their obvious limitations, we have learnt a lot from spectroscopic observations (maxwell_swarup; classen2002association; mann_formation; gopalswamy2005type, e.g.,). However, due to the large and potentially simultaneous variations in temporal and spectral structures, brightness temperatures and polarization properties of the solar radio bursts, spectro-polarimetric snapshot imaging with high dynamic range is needed to study these events in their full glory. Generations of former and presently operational radioheliographs like Culgoora Wild1967, Clark Lake Gregley1985, NoRH nakajima; takano2007upgrade; shibasaki2013long, the Nançay Radioheliograph kerdraon; saint2012decade and Gauribidnur Ramesh1998 etc. enabled routine imaging of the Sun at a few discrete frequencies, helping in the study of solar flares and noise storms. Despite this, the instrumentation, algorithmic and computational challenges posed by the needs of solar radio imaging have held back progress. More recent solar-dedicated instruments such as the Siberan Radioheliograph srh_2020, Mingantu Spectral Redioheliograph yan2021mingantu, Expanded Owens Valley Solar Array (gary2018microwave, EOVSA;), Owens Valley Radio Observatory - Long Wavelength Array (chhabra2021imaging, OVRO-LWA;), and Daocheng Solar Radio Telescope (Yan2023-DSRT, DSRT;), have made progress in addressing some of the limitations of earlier instruments (gary2023new, see review by). Alongside, over the last decade or so, a new generation of radio interferometers have emerged, e.g. the LOFAR lofar_paper, the MWA tingay2013murchison, MeerKAT meerkat_2016. In terms of their raw capabilities, these instruments tend to be best-in-class along some/many of the parameters of interest (e.g. uv-coverage, resolution, spectroscopic imaging capability, calibration fidelity, etc.) and are much better suited to the demanding requirements of solar imaging. These sophisticated instruments, all precursors or pathfinders for the Square Kilometre Array Observatory (SKAO), are designed to be versatile, capable of addressing a vast variety of science areas and optimized for synthesis observations of faint sources. Considerable effort is required and has been put in to adapt these instruments and the analysis procedures for solar observations (Kansabanik2022c; Kansabanik_2022; Kansabanik_2023; dey2025automated; kansabanik2025solar, e.g., ). These efforts have and continue to deliver considerable science returns (morosan_2014; reid_kontar; kansabanik2024spectropolarimetric; oberoi2023preparing; shilpi_type2; kansabanik2025solar; dey2025first, e.g.,).
@@
Being versatile and state-of-the-art, the observing time at these new generation facilities is highly oversubsribed. On the other hand, the solar radio bursts are infrequent, transient, and most importantly, unpredictable. Hence, blind scheduled solar observations are not only an extremely inefficient way to capture these bursts in large numbers, they also lead to a high likelihood of missing big solar events.
@@
One way to overcome this observational limitation is by optimally utilizing the available observing time using a robust and reliable automated near-real-time triggered observation framework. By enabling one to use precious observing time only when some solar activity is known to have just taken place, such a system can vastly increase the efficiency of the limited available observing time to capture instances of solar activity. Having such systems will not only help in boosting efficiency of Solar research with the existing instruments but also will pave a way to prepare for optimized solar observations with the upcoming SKAO.
@@
In radio astronomy, triggered observations have been widely used to capture astrophysical transient activities such as FRBs, GRB afterglows, or pulsar glitches where the response times of seconds to hours are acceptable (frb_trigger_2020; grb_triggered_2024, e.g.,). Solar radio burst triggers need near-real time response and low-latency setups to configure the interferometer to observe the Sun as soon as the trigger is received. We have developed such a framework for the low-frequency precursor of the SKAO, the MWA, based on near-real-time data from the Yamagawa solar radio spectrograph Iwai_2017. We have named this framework Solar Triggered Observations of Radio bursts using MWA and Yamagawa (STORMY).
@@
The paper is organized as follows. An overview of the pipeline is given in section 2. Section 3 and 4 describe the preprocessing and the event detection, respectively. Trigger generation and observation are shown in section 5. We present the performance test of STORMY in section 6. Section 7 shows a few events captured by our pipeline. Finally, we discuss the application of this algorithm to other facilities in section 8.
@@ 2 Overview of Real-time Burst Detection and Triggering Pipeline
While there have been some attempts to develop automated algorithms for the detection of solar radio bursts from dynamic spectrum, they tend to be focused on the detection and characterization of the type of bursts (e.g. type IIs or IIIs). Earlier works by Lobzinn and bonnin2011automated proposed the use of the Hough transform hough1962 on the dynamic spectra. Dayal have developed an automated burst detection pipeline using e-CALLISTO benz2005callisto; benz2009world data, and salmane have developed a Constant False Alarm Rate (CFAR) based technique for automatic detection of radio bursts in data from the Nançay Decametric Array boischot1980new. In recent times, scientists have also successfully implemented different convolutional neural network (CNN) based deep learning models to detect and categorize solar radio bursts (scully2021type; orue2023automatic; deng2024real, e.g.,). However, these algorithms were developed for searching and classifying radio bursts present in the archival data, where the entire morphological information of the burst is available in the dynamic spectrum.
In case of generating a trigger for scheduling an interferometric observation, one does not have the luxury of allowing the full solar radio event to play out before raising the trigger. Developing such a framework to conduct robust and efficient triggered observations requires the following capabilities: (1) near-real-time access to appropriate data from solar-dedicated telescopes, based on which a trigger can be generated; (2) algorithms and software tools needed to ingest and process these data, including reliably rejecting any radio frequency interference (RFI), to provide a trigger; (3) the framework required for this trigger to be communicated to the telescope under consideration and, (4) scheduling the observations and the acquisition of data. Here we focus on the first two of the aspects mentioned here, and the latter two have already been implemented at the MWA for other astronomical observations and have been adapted for the present purpose.
@@ 2.1 Yamagawa and MWA
Our solar observing trigger is derived from the data received from the Yamagawa heliospectrograph, located in Japan at a latitude and longitude of 31.204N and 130.617E, respectively. It observes the Sun daily from rise to set in the frequency band from 70 MHz to 9 GHz. Since Yamagawa is at a similar longitude as the MWA (26.703S, 116.671E) and the MWA only observes the Sun above 30 elevation, Yamagawa has complete overlap with the solar observing window of the MWA.
Figure 1: This figure shows an example of the real-time update of data from Yamagawa spectrograph and the MWA voltage buffer. For example: the chunk a is available at 2:46 and the next chunk b is available at 2:48. When a trigger is sent, the buffer data (red block) is saved and a regular correlator observation (blue arrow) is triggered. We lose about 70–80 s of data even after recovering 160 s of data from the MWA buffer.
We have optimized the Yamagawa system to enable a custom near-real-time datastream. This datastream provides dynamic spectrum at 2-minute intervals with a latency of 2 minutes and is better suited to our needs than the publicly available data. The data have a spectral and temporal resolution of 1 MHz and 1 second, respectively.
@@
The MWA observes in the 80–300 MHz band and offers 30.72 MHz of instantaneous processed bandwidth. This instantaneous bandwidth is available in 24 chunks of 1.28 MHz each, which can be distributed flexibly across the entire observing frequency range. Solar observations are often carried in the so called picket fence mode where the observing bandwidth is distributed in 1.28 MHz chunks across the full 80–300 MHz band in a log-spaced manner.
The MWA also offers a unique ability to store the voltage data for each of the polarizations from each of the tiles in a ring buffer morrison2023mwax. These data are extremely voluminous ($\sim$56 TB/hour for 128 tiles) and with 128 tiles the ring buffer can store upto 160s160\ $\mathrm${s} of raw voltage data. The voltage buffer offers two key advantages – first, it provides access to the data for up to 160s160\ $\mathrm${s} prior to the trigger arriving at the MWA and thus allowing us to recover much of the event data lost due to the latency in raising the trigger; and second, it provides flexibility in the time and frequency resolution with which to compute correlations (visibilities) down to the Nyquist sampling rate. On command, this buffer data can be dumped to disk and can later be passed through an offline correlator to obtain visibility data and reduce data volume. The buffer is filled by the datastream from the pointing of the last observation. Hence, we configure the MWA to point at the Sun in the picket fence mode for all of the unallocated daytime hours, when the Sun is above the elevation limit of the MWA (30). So even though no visibility data is recorded, this ensures that the buffer is populated with solar data. Figure 1 shows how the near real-time datastream comes from the Yamagawa spectrograph, the time taken to raise the trigger and the start of observations at the MWA. It is clear from the diagram that due to the buffer available at the MWA (red block), we lose a very small amount of data (blue block) due to the latency of the Yamagawa data.
@@ 2.2 Structure of the Pipeline
For an overview of the full pipeline, see figure 2. After the near-real-time data is downloaded, the data flows through the following parts:
@@ 3.1 Bandpass Correction
The bandshape of an instrument is its spectral response and can vary significantly across the observing band. The observed intensity of the signal that is recorded by each spectral channel gets modified by the instrumental bandshape. Hence, we need to correct for the power levels of each spectral channel by the instrumental bandshape. Although the same-day data can also be used for band shape calculation, identifying quiet time without any solar bursts in the real-time data becomes computationally inefficient, and also, if a very long duration solar signal is present in the data, the median might get affected. Hence, to calculate the bandshape of the Yamagawa spectrograph, we use previous full-day data ($\sim$8 hours). This data is split into chunks each spanning 180s180\ $\mathrm${s}. Now, we calculate the bandshape of each chunk ($B_{i}(\nu)$) by taking median along the temporal axis and then we calculate the final bandshape by taking median of all chunks ($B(\nu)=Median[B_{i}(\nu)]$). This median bandshape is used to normalize the dynamic spectra as follows, $S(\nu,t)=\frac{I(\nu,t)}{B(\nu)}$, where, $I(\nu,t)$ denotes the raw dynamic spectrum. This normalization removes instrumental spectral response and solar features become more easily detectable. Figure 3 shows a sample raw data and the previous day bandshape used for bandpass correction. We will use this data throughout this paper to show further steps.
Figure 4: This figure shows the bandpass normalized dynamic spectrum where the varying signals are easily detectable. The yellow rectangles show a periodic RFI that is seen every 300 s. The red arrows show narrow band persistent RFIs.
@@ 3.2 RFI Mitigation
After the bandpass normalization, different spectrally and temporally varying features can be seen with enhanced contrast. However, the dynamic spectrum has different types of unwanted contamination from RFI which can sometimes mimic the true solar signal. We need to remove them before starting to detect the solar signal. Otherwise, we may get false positive detections. Analyzing the archival Yamagawa data, we identified a few different types of RFI that can be present in the data (Figure 4) – (1) periodic RFI, (2) narrow band and persistent RFI and (3) transient RFI.
First, we remove the periodic RFI. They are present every day and have a periodicity of five minutes with a duty-cycle of $\sim 10\ \mathrm{s}$, but are fragmented across the channels (Refer to the bottom left panel of figure 5). Although they are fragmented, they appear exactly after 300 s in each channel (top left panel of figure 5). Let us denote $S_{\nu}(t)$ as time series a channel. We use the find_peaks module from scipy to find peaks with periodicity of 300 s by using the following parameters:
@@
Figure 5: This figure shows how the periodic RFI is cleaned. The top left panel shows the dynamic spectrum with fragmented periodic RFI. The top right panel shows the dynamic spectrum after cleaning these RFI. Bottom panels: Zoomed-in view of the interval between the red lines, showing before (left) and after (right) cleaning the RFI.
Next, we clean any persistent narrow-band RFI. They span only a few channels in frequency, deviating largely from the background and persistently present for a long time. Since solar emissions can also have significant deviations from the background, extra care should be taken while filtering these RFI. We used a median filter with a sufficiently small moving window (5 MHz) to arrive at a smoothed representation of the spectrum at a timestamp, $t_{0}$. This is denoted by $S_{\mathrm{m}}(\nu,t=t_{0})$. We calculate the modified $z$-score of the quantity $\delta(\nu,t=t_{0})=[S(\nu,t=t_{0})-S_{m}(\nu,t=t_{0})]$. Compared to regular z-score calculation, we use the modified z-score since it uses median absolute deviation ($MAD(x)=Median(|x_{i}-Median(x)|)$) and is less affected by extreme outliers. In our case it is calculated as follows:
@@
$S^{\prime}(\nu,t)=\begin{cases}S(\nu,t),&\text{if }z_{m}<10\\ S_{m}(\nu,t),&\text{otherwise}\end{cases}$ (3)
@@
Figure 6: The top left and right panels show the dynamic spectrum before and after narrow band RFI removal, respectively. The bottom left and right panels show the same for spectra at UTC 03:44, as shown by the red vertical line on the dynamic spectrum in the top panel. The red trace in bottom left panel shows the median smoothed spectra ($S_{m}(\nu)$) of the same.